### 01 Apr 2026 v1

# Rigidity of the Borell-Brascamp-Lieb Inequality on Weighted Riemannian Manifolds

Abstract
In this paper we discuss some results regarding the rigidity of the Borell-Brascamp-Lieb inequality and the Brunn-Minkowski inequality. We show a theorem of rigidity on curvature and measure of the Borell-Brascamp-Lieb inequality, a generalisation of the curvature rigidity theorem by Balogh-Kristály in [2] to the weighted setting. We present some rigidity results of the Brunn-Minkowski inequality and a few further open problems.
1 Introduction
Since the seminal works on the application of optimal transport in geometric analysis including [19] by Otto-Villani, [20] by Sturm and [13] by Lott-Villani, it became possible to study a wide range of analytical problems on Riemannian manifolds using the tool of optimal transport. In particular, the theory of optimal transport has been a powerful tool in the study of functional and geometrical inequalities on Riemannian manifolds with curvature and dimension bounds.
First proved in 1952 by Henstock and MacBeth on $\mathbb{R}$ in [12] and later developed by Borell and Brascamp-Lieb in [3] and [4], the Borell-Brascamp-Lieb inequality is an integral inequality concerned with the $L^{1}$-norm of two functions and that of the mean of them.
On the $n$-dimensional Euclidean space $\mathbb{R}^{n}$, the Borell-Brascamp-Lieb inequality states that if functions $f,g,h>0$ such that for $s\in(0,1)$,
$h((1-s)x+sy)\geq\mathcal{M}^{p}_{s}(f(x),g(y))\quad\forall(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n},$
then
$\|h\|_{L^{1}}\geq\mathcal{M}^{\frac{p}{1+np}}_{s}(\|f\|_{L^{1}},\|g\|_{L^{1}})$
where for $p\in[-\frac{1}{n},\infty]$ and real numbers $a,b\geq 0$,
[…]
is the $p$-mean, with the conventions $\mathcal{M}^{\infty}_{s}(a,b)=\max\{a,b\}$, $\mathcal{M}^{0}_{s}(a,b)=a^{1-s}b^{s}$ and $\mathcal{M}^{-\infty}_{s}=\min\{a,b\}$.
The Borell-Brascamp-Lieb inequality was proved on Riemannian manifolds by Cordero-Erausquin, McCann and Schmuckenschläger in [6]. It was then generalised further to metric measure spaces by Bacher in [1]. The Borell-Brascamp-Lieb inequality implies a number of functional and geometric inequalities including the Brunn-Minkowski inequality and the Prékopa-Leindler inequality, a diagram showing their relationship can be seen in [11].
Rigidity of the Borell-Brascamp-Lieb inequality on $\mathbb{R}^{n}$ has been studied since 1970s, various papers including [8] by Dubuc since then have discussed this problem. Rigidity of it on (unweighted) Riemannian manifolds with curvature bounded from below was shown by Balogh-Kristály in [2].
This paper generalises the curvature rigidity result of the Borell-Brascamp-Lieb inequality on Riemannian manifolds by Balogh and Kristály, Theorem 4.1 of [2], to weighted Riemannian manifolds. The main result of this paper states that the rigidity of the Borell-Brascamp-Lieb inequality implies identical sectional curvatures and certain behaviour of the measure.
Main Theorem (Theorem 2).
On a weighted Riemannian manifold $(M,g,m=e^{-\psi}\mathrm{vol}_{g})$, satisfying $\mathrm{Ric}_{m,N}\geq Kg$ where $n\leq N<\infty$ and $K\in\mathbb{R}$, equality of the Borell-Brascamp-Lieb inequality implies that the $N$-Ricci curvature along the geodesic of transport $\gamma(t):=\mathcal{F}_{t}(x)$ equals $K$ and the sectional curvatures along the geodesic $\gamma$ are identically $\frac{K}{N-1}$. The measure behaves as an (Nn)(N-n)-th power of a linear combination of trignometric functions along the geodesic $\gamma(t)$ when $K>0$; as an (Nn)(N-n)-degree polynomial of $t$ along $\gamma(t)$ when $K=0$; and as an (Nn)(N-n)-th power of of a linear combination of hyperbolic functions along $\gamma(t)$ when $K<0$ for some constants $c>0$.
The article is structured as follows. In section 2, we first demonstrate a proof of an interpolation inequality on weighted Riemannian manifolds, followed by a proof of the Borell-Brascamp-Lieb inequality on weighted Riemannian manifolds. In section 3, we discuss the rigidity of the Borell-Brascamp-Lieb inequality and prove the main theorem, Theorem 2, and in section 4 we discuss the Brunn-Minkowski inequality and its rigidity, as a consequence of the rigidity of the Borell-Brascamp-Lieb inequality. At the end in section 5 we present a few related open questions.
Acknowledgements
I would like to express my deep gratitude to Prof. Shin-ichi Ohta for his guidance, suggestions and discussions. I would also like to thank Dr. Hiroshi Tsuji and Mr. Zisu Zhao for some helpful discussions.
This work was supported by JST SPRING, Grant Number JPMJSP2138.
2 Preliminaries
Throughout the paper, we consider a smooth weighted Riemannian manifold $(M,g,m)$ of dimension $n$, where $m=e^{-\psi}\mathrm{vol}_{g}$, satisfying the condition $\mathrm{Ric}_{m,N}(\cdot)\geq Kg(\cdot,\cdot)$, $N\in[n,\infty)$ and $K\in\mathbb{R}$. Recall that the $N$-Ricci curvature $\mathrm{Ric}_{m,N}$ is defined to be:
$\mathrm{Ric}_{m,N}(v):=\left\{\begin{array}[]{cl}\mathrm{Ric}_{m}(v)-\frac{(\partial_{v}\psi)^{2}}{N-n}\quad&\mbox{for $N\in(n,\infty)$},\\ \mathrm{Ric}_{m}(v)\quad&\mbox{for $N=\infty$},\\ \mathrm{Ric}(v)\quad&\mbox{for $N=n$, $\psi=\mathrm{const.}$}\end{array}\right$.
where $\mathrm{Ric}_{m}(v)=\mathrm{Ric}(v)+\mathrm{Hess}_{\psi}(v,v)$.
We denote $\mathcal{P}(M)$ to be the space of Borel probability measures on $M$, and let $\mathcal{P}^{2}(M)$ be the space of measures with finite second moment
$\mathcal{P}^{2}(M)=\Big\{\mu\in\mathcal{P}(M)\Big|\int_{M}d(x,x_{0})^{2}d\mu<\infty\text{ for arbitrary }x_{0}\in M\Big\}.$
By the theorem of Brenier-McCann, given two measures on $(M,g,m)$, $\mu_{0}=\rho_{0}m=\rho_{0}e^{-\psi}\mathrm{vol}_{g}$ and $\mu_{1}=\rho_{1}m=\rho_{1}e^{-\psi}\mathrm{vol}_{g}$ in $\mathcal{P}^{2}(M)$ that are absolutely continuous with respect to $m$, there is a unique optimal transference plan from $\mu_{0}$ to $\mu_{1}$, given by $\mathcal{F}_{t}$, $\mu_{t}=(\mathcal{F}_{t})_{\#}\mu_{0}$, the push-forward of $\mu_{0}$ by $\mathcal{F}_{t}(x):=\exp_{x}(t\nabla f(x))$, and $f:M\rightarrow\mathbb{R}$ is a $\frac{d^{2}}{2}$-convex function. We refer to [21] for the fundamentals of optimal transport theory, see also [15] and [10] for the theory of optimal transport on Riemannian manifolds.
We consider the weighted Jacobian $J^{\psi}_{t}(x):=e^{\psi(x)-\psi(\mathcal{F}_{t}(x))}\|(D\mathcal{F}_{t}(x))\|$ along the geodesic $\gamma(t):=\mathcal{F}_{t}(x)$. We have the following concavity condition on the weighted Jacobian:
$J^{\psi}_{t}(x)^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}$ (3)
for $\mu_{0}$-a.e. $x$ such that $\mathcal{F}_{t}$ is differentiable at $x$ and all $t\in[0,1]$, with the function $\beta^{t}_{K,N}$ being defined as:
$\beta^{t}_{K,N}(r)=\Big(\frac{s_{K,N}(tr)}{ts_{K,N}(r)}\Big)^{N-1}$
where
$s_{K,N}(r):=\left\{\begin{array}[]{cl}\sqrt{\frac{N-1}{K}}\sin\Big(r\sqrt{\frac{K}{N-1}}\Big)\quad&\mbox{if $K>0$},\\ r\quad&\mbox{if $K=0$},\\ \sqrt{-\frac{N-1}{K}}\sinh\Big(r\sqrt{\frac{-K}{N-1}}\Big)\quad&\mbox{if $K<0$}.\end{array}\right$.
It is known that the density function $\rho_{t}$ given by $\mu_{t}=\rho_{t}m$ satisfies the Monge-Ampère equation
$\rho_{0}(x)=\rho_{t}(\mathcal{F}_{t}(x))J^{\psi}_{t}(x)\text{ for }\mu_{0}\text{-a.e. }x.$ (5)
The interpolation inequality (3) on unweighted Riemannian manifolds was shown in the proof of Theorem 1.7 of [20], and its Finsler version was shown in the proof of Theorem 1.2 of [18]. Here we present a proof on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$, along the computation shown in the proof of Theorem 1.7 of [20], with an addition of the factor due to measure to the proof in [20], for later use. See also pages 393-401 of [22].
Lemma 1.
On a weighted Riemannian manifold $(M,g,m=e^{-\psi(x)}\mathrm{vol}_{g})$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$, the weighted Jacobian $J^{\psi}_{t}(x):=e^{\psi(x)-\psi(\mathcal{F}_{t}(x))}\|(D\mathcal{F}_{t}(x))\|$ defined along the geodesic $\gamma(t):=\mathcal{F}_{t}(x)$, where $\mathcal{F}_{t}(x)=\exp_{x}(t\nabla f(x))$ described as above satisfies the concavity condition
$J^{\psi}_{t}(x)^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}.$
Proof.
We fix $x\in\mathrm{supp}\mu_{0}$ on $M$, and consider the geodesic
$\gamma(t):[0,1]\rightarrow M,t\mapsto\mathcal{F}_{t}(x)$
representing the path of transport starting from $x$.
When $\gamma(t)$ is constant, inequality (3) is trivial as $J_{t}^{\psi}(x)=1$ identically.
When $\gamma(t)$ is nonconstant, at $x$, we consider the tangent space $T_{x}M=T_{\mathcal{F}_{0}(x)}M$ and an orthonormal frame $\{e_{0}^{i}\}_{i=1}^{n}$, where we denote $e^{1}_{0}$ to be the unit tangent of the geodesic: $\frac{\gamma^{\prime}(0)}{\|\gamma^{\prime}(0)\|}$. We can thus define n1n-1 Jacobi fields $J_{i}(t)=\frac{\partial}{\partial s}\exp_{x}(t(\nabla f(x)+se^{i}_{0}))$ for $i\in\{2,...,n\}$ that are orthogonal to $\gamma^{\prime}(t)$, and the vectors $\{J_{i}(t)\}$ form a basis of $\{\gamma^{\prime}(t)\}^{\perp}\subset T_{\mathcal{F}_{t}(x)}M$ along the entire geodesic $\gamma$.
We denote $A_{t}(x)$ to be the differential map:
$A_{t}(x):=D\mathcal{F}_{t}(x):T_{x}M\rightarrow T_{\mathcal{F}_{t}(x)}M.$
It is known that $A_{t}$ satisfies the Riccati equation
$D_{t}D_{t}A_{t}(x)+R(A_{t}(x),\gamma^{\prime}(t))\gamma^{\prime}(t)=0$
with the initial condition $A_{0}=\mathrm{Id}$ and $D_{t}A_{t}|_{t=0}=-\mathrm{Hess}\psi$, see page 142 of [20].
Since $\{J_{i}\}_{i\geq 2}$ form a basis of the vector space $\{\gamma^{\prime}(t)\}^{\perp}$, and since for $i\geq 2$, $\langle J_{i}^{\prime}(t),\gamma^{\prime}\rangle=0$ by Gauß lemma, we can define, as in [20], the matrix $U(t)$ to be such that $U(t)A_{t}:=D_{t}A_{t}$. And the Riccati equation can also be written as
$U^{\prime}+U^{2}+RA^{-1}=0\text{, or }U^{\prime}+U^{2}+R(\cdot,\partial_{t}\mathcal{F}_{t})\partial_{t}\mathcal{F}_{t}=0.$
At $\gamma(t)$ we set an orthonormal basis $\{e^{i}_{t}\}$ with $e^{1}_{t}:=\gamma^{\prime}(t)/\|\gamma^{\prime}(t)\|$ being the tangent vector of $\gamma(t)$ of unit length. In terms of this basis, we set $U_{ij}(t)=g(e^{i}_{t},U(t)e^{j}_{t})$, $\lambda_{t}=1+\int_{0}^{t}U_{11}(s)ds$, and $f(t)=\exp(\lambda_{t})$, the Riccati equation implies that
$\frac{d}{dt}U_{11}(t)+\sum_{j=1}^{n}U_{1j}^{2}=0\text{, therefore }-\lambda^{\prime\prime}=-\frac{d}{dt}U_{11}(t)=\sum_{j=1}^{n}U_{1j}^{2}\geq U_{11}^{2}=(\lambda^{\prime})^{2}$
which implies further that:
$f^{\prime\prime}=\lambda^{\prime\prime}f+(\lambda^{\prime})^{2}f\leq 0.$
We let $y_{t}:=\mathrm{log}J_{t}$ where $J_{t}:=\mathrm{det}A_{t}=\|D\mathcal{F}_{t}(x)\|$, $y^{\psi}_{t}:=\mathrm{log}J^{\psi}_{t}=\psi(x)-\psi\circ\mathcal{F}_{t}(x)+y_{t}$ and $\alpha_{t}:=y^{\psi}_{t}-\lambda_{t}$. Define
$u(t):=e^{\alpha_{t}}=\frac{J^{\psi}_{t}}{f(t)},$
since $y^{\prime}=\mathrm{tr}U$, $y^{\prime\prime}=\mathrm{tr}(U^{\prime})$, we have
$\alpha^{\prime\prime}=y^{\prime\prime}-\lambda_{t}^{\prime\prime}-(\psi\circ\mathcal{F}_{t})^{\prime\prime}$ $=$ $-\mathrm{Ric}(\gamma^{\prime})-\mathrm{tr}(U^{2})-U_{11}^{\prime}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$=$ $-\mathrm{Ric}(\gamma^{\prime})-\sum_{i,j=1}^{n}U_{ij}^{2}+\sum_{j=1}^{n}U_{1j}^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\sum_{i,j=2}^{n}U_{ij}^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\frac{1}{n-1}\Big(\sum_{i=2}^{n}U_{ii}\Big)^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime}).$
By the Cauchy-Schwarz inequality, it follows that
$\alpha^{\prime\prime}=y^{\prime\prime}-\lambda_{t}^{\prime\prime}-(\psi\circ\mathcal{F}_{t})^{\prime\prime}$ $\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\frac{(y^{\prime}-\lambda^{\prime})^{2}}{n-1}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$ (6)
$=$ $-\mathrm{Ric}_{m,N}(\gamma^{\prime})-\frac{(\partial_{\gamma^{\prime}}\psi)^{2}}{N-n}-\frac{(y^{\prime}-\lambda^{\prime})^{2}}{n-1}$
$\leq$ $-\mathrm{Ric}_{m,N}(\gamma^{\prime})-\frac{(\alpha^{\prime})^{2}}{N-1}.$
Inequality (6) and the hypothesis $\mathrm{Ric}_{m,N}\geq K$ imply that $u$ satisfies the differential inequality, as in pages 143-144 of [20],
$(u^{\frac{1}{N-1}})^{\prime\prime}\leq-\frac{\mathrm{Ric}_{m,N}(\gamma^{\prime})}{N-1}u^{\frac{1}{N-1}}\leq-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}}\text{, where }r=d(x,\mathcal{F}_{1}(x))$
which implies that $u^{\frac{1}{N-1}}(t)$ satisfies the concavity condition
$u^{\frac{1}{N-1}}(t)\geq\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1).$
After multiplying $f^{\frac{1}{N}}$ and $u^{\frac{1}{N}}$ and applying the Hölder’s inequality, we have
$(J^{\psi}_{t})^{\frac{1}{N}}$ $=$ $(f(t)u(t))^{\frac{1}{N}}=f^{\frac{1}{N}}(t)(u^{\frac{1}{N-1}}(t))^{\frac{N-1}{N}}$
$\geq$ $((1-t)f(0)+tf(1))^{\frac{1}{N}}\Big(\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)\Big)^{\frac{N-1}{N}}$
$\geq$ $(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}$
showing the inequality (3). ∎
Along [1], we can show a version of Borell-Brascamp-Lieb inequality on weighted Riemannian manifolds:
Theorem 1.
Let $f_{0},f_{1},h\geq 0$ be functions on a weighted Riemannian manifold $(M,g,m)$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$. Assume that for Borel subsets $A,B\subset M$, $\int_{A}f_{0}dm=\int_{B}f_{1}dm=1$ and
$h(z)^{-\frac{1}{N}}\leq(1-t)\beta^{1-t}_{K,N}(d(x,y))^{\frac{1}{N}}f_{0}(x)^{-\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,y))^{\frac{1}{N}}f_{1}(x)^{-\frac{1}{N}}$
for all $(x,y)\in A\times B,z\in Z_{t}(x,y)$, where
$Z_{t}(x,y):=\{z\in M|d(x,z)=td(x,y),d(z,y)=(1-t)d(x,y)\}$
is the set of $t$-intermediate points between $x$ and $y$. Then $\int_{M}hdm\geq 1$.
The proof of the above theorem is essentially the same as the proof of the main theorem of [1], which is similar to the proof of the main theorem of [6] by replacing the volume distortion coefficient by the function $\beta^{t}_{K,N}$.
Proof.
By the assumption that $\mathrm{Ric}_{m,N}\geq Kg$, we have the concavity condition of the weighted Jacobian (3).
We let $\rho_{0}:=f_{0}|_{A}$, $\rho_{1}:=f_{1}|_{B}$. There exists an optimal transference plan with density $\rho_{t}(x)$, satisfying Monge-Ampère equation (5), from $\mu_{0}=\rho_{0}m$ to $\mu_{1}=\rho_{1}m$ and induced by a map $\mathcal{F}_{t}$ such that
$J^{\psi}_{t}(x)\cdot\rho_{t}\circ\mathcal{F}_{t}(x)=\rho_{0}(x)=f_{0}(x)\text{ for }m\text{-a.e. }x\in A,$
and we have, letting $y=\mathcal{F}_{1}(x),$
$\Big(\frac{f_{0}(x)}{\rho_{t}(\mathcal{F}_{t}(x))}\Big)^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}(d(x,y))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,y))^{\frac{1}{N}}\Big(\frac{f_{0}(x)}{f_{1}(\mathcal{F}_{1}(x))}\Big)^{\frac{1}{N}},$
$\Big(\frac{1}{\rho_{t}(\mathcal{F}_{t}(x))}\Big)^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}(d(x,y))^{\frac{1}{N}}\frac{1}{f_{0}(x)^{\frac{1}{N}}}+t\beta^{t}_{K,N}(d(x,y))^{\frac{1}{N}}\frac{1}{f_{1}(\mathcal{F}_{1}(x))^{\frac{1}{N}}}\geq h(\mathcal{F}_{t}(x))^{-\frac{1}{N}},$
hence
$h\circ\mathcal{F}_{t}\geq\rho_{t}\circ\mathcal{F}_{t}.$
Thus, we conclude
$\int_{M}hdm$ $\geq$ $\int_{\mathcal{F}_{t}(A)}hdm\geq\int_{M}\mathbf{1}_{\mathcal{F}_{t}(A)}\rho_{t}dm$
$=\int_{A}\mathbf{1}_{\mathcal{F}_{t}(A)}\circ\mathcal{F}_{t}(x)\cdot\rho_{t}\circ\mathcal{F}_{t}(x)J^{\psi}_{t}dm=\int_{A}\mathbf{1}_{\mathcal{F}_{t}(A)}(\mathcal{F}_{t}(x))d\mu_{0}$
$=1.$
The above theorem can be expressed in terms of the $p$-mean via the same method of scaling as the main theorem of [6] was generalised to Corollary 1.1 of [6].
Corollary 1.
Let $f_{0},f_{1},h\geq 0$ be functions on a weighted Riemannian manifold $(M,g,m)$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$. Assume that
$h(z)\geq\mathcal{M}^{p}_{t}\Big(\frac{f_{0}(x)}{\beta^{1-t}_{K,N}(d(x,y))},\frac{f_{1}(y)}{\beta^{t}_{K,N}(d(x,y))}\Big)\text{ }\forall(x,y)\in M\times M,z\in Z_{t}(x,y)$
then
$\int_{M}hdm\geq\mathcal{M}_{t}^{\frac{p}{1+Np}}\Big(\int_{M}f_{0}dm,\int_{M}f_{1}dm\Big)$
where
[…]
is the $p$-mean, for $p\in\big[-\frac{1}{N},\infty\big]$, as introduced in the introduction.
3 Rigidity of the Borell-Brascamp-Lieb inequality
Suppose that for some $t_{0}\in(0,1)$, the Borell-Brascamp-Lieb inequality on a weighted Riemannian manifold as in Theorem 1 attains equality.
We therefore have, with the constructions above, for $\mu_{0}$-a.e. $x\in A$,
$\int_{M}hdm$ $=$ $\int_{\mathcal{F}_{t_{0}}(A)}hdm=\int_{M}\mathbf{1}_{\mathcal{F}_{t_{0}}(A)}\rho_{t_{0}}dm$
$=$ $\int_{A}\mathbf{1}_{\mathcal{F}_{t_{0}}(A)}\circ\mathcal{F}_{t_{0}}(x)\cdot\rho_{t_{0}}\circ\mathcal{F}_{t_{0}}(x)J^{\psi}_{t_{0}}dm=\int_{A}\mathbf{1}_{\mathcal{F}_{t_{0}}(A)}(\mathcal{F}_{t_{0}}(x))d\mu_{0}$
$=$ $1$
implying that on $A$, $h\circ\mathcal{F}_{t_{0}}=\rho_{t_{0}}\circ\mathcal{F}_{t_{0}}$ almost everywhere and hence
$J^{\psi}_{t_{0}}(x)^{\frac{1}{N}}=(1-t_{0})\beta^{1-t_{0}}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t_{0}\beta^{t_{0}}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}.$ (8)
By the comparison of differential inequalities, the interpolation inequality (8) attaining equality at one value of $t$ implies that the inequality attains equality for all $t\in[0,1]$, as follows.
Lemma 2.
If the weighted Jacobian $J^{\psi}_{t}$ satisfying (3) attains equality at some $t_{0}\in(0,1)$, then equality of the inequality (3) holds for all $t\in[0,1]$.
Proof.
The result is shown by comparison of ODE. We give a proof for $K>0$ and the other situations can be handled in the same way. Recall that with the definition of $f$ and $u$ as in the proof of Lemma 1, the inequality (3) is derived by:
$(J^{\psi}_{t})^{\frac{1}{N}}$ $=$ $(fu)^{\frac{1}{N}}=f^{\frac{1}{N}}(u^{\frac{1}{N-1}})^{\frac{N-1}{N}}$
$\geq$ $((1-t)f(0)+tf(1))^{\frac{1}{N}}\Big(\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)\Big)^{\frac{N-1}{N}}$
$\geq$ $(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}.$
Claim 1: If equality holds at $t_{0}$, then $f(t_{0})=(1-t_{0})f(0)+t_{0}f(1)$ and $u(t_{0})^{\frac{1}{N-1}}=\frac{s_{K,N}((1-t_{0})r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(t_{0}r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$ where $r=d(x,\mathcal{F}_{1}(x))$.
Proof of Claim 1: Suppose, on the contrary, that the claim is not true, that
$f(t_{0})>(1-t_{0})f(0)+t_{0}f(1)$
then we will have that $u(t_{0})^{\frac{1}{N-1}}<\frac{s_{K,N}((1-t_{0})r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(t_{0}r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$, which is a contradiction, as the differential inequality
$(u^{\frac{1}{N-1}})^{\prime\prime}\leq-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}}$
implies otherwise.
Or similarily, if $u(t_{0})^{\frac{1}{N-1}}>\frac{s_{K,N}((1-t_{0})r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(t_{0}r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$, we will get the same contradiction for $f$.
Claim 2: If $u$ satisfies $(u^{\frac{1}{N-1}})^{\prime\prime}\leq-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}}$ for $t\in(0,1)$ and there exists $t_{0}\in(0,1)$ such that $u(t_{0})^{\frac{1}{N-1}}=\frac{s_{K,N}((1-t_{0})r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(t_{0}r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$, then
$u^{\frac{1}{N-1}}(t)=\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)\quad\forall t\in[0,1].$
Proof of Claim 2: We define $w(t):=\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$. Note that
$w^{\prime\prime}(t)=-\frac{Kr^{2}}{N-1}w.$
Thus, $u^{\frac{1}{N-1}}-w$ satisfies the differential inequality
$(u^{\frac{1}{N-1}}-w)^{\prime\prime}\leq\frac{Kr^{2}}{N-1}(w-u^{\frac{1}{N-1}})\leq 0$
and has boundary values $u^{\frac{1}{N-1}}(0)-w(0)=u^{\frac{1}{N-1}}(1)-w(1)=0$. Since also at $t_{0}$, $u^{\frac{1}{N-1}}-w=0$, by the maximum principle of subharmonic functions, we conclude
$u^{\frac{1}{N-1}}-w=0\quad\forall t\in[0,1].$
The lemma follows from Claim 2. ∎
We can thus show the main result by concluding:
Theorem 2.
Rigidity of the Borell-Brascamp-Lieb inequality on weighted Riemannian manifold $(M,g,m)$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$ implies, for a.e. $x\in\mathrm{supp}f_{0}$, letting $\gamma(t):=\mathcal{F}_{t}(x)$, the following hold:
  1. 1.
    For all $t\in(0,1)$, the $N$-Ricci curvature $\mathrm{Ric}_{m,N}(\gamma^{\prime}(t))$ equals $Kg(\gamma^{\prime},\gamma^{\prime})$, and the sectional curvatures along the geodesic $\gamma$ are identically $\frac{K}{N-1}$.
  2. 2.
    The measure behaves as an (Nn)(N-n)-th power of a linear combination of trignometric functions along the geodesic $\gamma(t)$ when $K>0$; as an (Nn)(N-n)-degree polynomial of $t$ along $\gamma(t)$ when $K=0$; and as an (Nn)(N-n)-th power of a linear combination of hyperbolic functions along $\gamma(t)$ when $K<0$, for some constants $c>0$.
Proof.
We assume $N>n$ in the proof, as the case of $N=n$ reduces to [2]. With the above lemma and the construction of the functions $\alpha,y,\lambda_{t},u,f$ and the definition of the matrix $U$ as in the proof of Lemma 1, we have for all $t\in(0,1)$,
$(J^{\psi}_{t})^{\frac{1}{N}}$ $=$ $(fu)^{\frac{1}{N}}=f^{\frac{1}{N}}(u^{\frac{1}{N-1}})^{\frac{N-1}{N}}$
$=$ $((1-t)f(0)+tf(1))^{\frac{1}{N}}\Big(\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)\Big)^{\frac{N-1}{N}}$
$=$ $(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}$
which implies that f(t)=(1t)f(0)+tf(1)f(t)=(1-t)f(0)+tf(1) and the function $u$ satisfies
$(u^{\frac{1}{N-1}})^{\prime\prime}$ $=$ $-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}}.$
Since we have the chain of differential inequality
$(u^{\frac{1}{N-1}})^{\prime\prime}\leq-\frac{\mathrm{Ric}_{m,N}(\gamma^{\prime})}{N-1}u^{\frac{1}{N-1}}\leq-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}},$
by squeezing, it can be concluded that along the geodesic $\gamma$, $\mathrm{Ric}_{m,N}(\gamma^{\prime})=Kg(\gamma^{\prime},\gamma^{\prime})$.
The differential inequality of $u$ attaining equality also implies that the differential inequality (6) attains equality:
$\alpha^{\prime\prime}=-\mathrm{Ric}_{m,N}(\gamma^{\prime})-\frac{(\alpha^{\prime})^{2}}{N-1}.$ (9)
The equality being attained has two implications. The first implication is that the matrix $V$, defined to be $V:=(U_{ij})_{i,j=2,...,n}$, satisfies the condition $\mathrm{tr}(V^{2})=\frac{1}{n-1}(\mathrm{tr}V)^{2}$, that it equals the scalar multiple of the identity matrix $V=\xi(t)\mathrm{Id}_{n-1}$, which further implies that the sectional curvatures along the geodesic $\gamma(t)$ are all identical by the Riccati equation.
The second implication is that the Cauchy-Schwarz inequality between the tuples
$\Big(-\frac{\partial_{\gamma^{\prime}}\psi}{\sqrt{N-n}},\frac{y^{\prime}-\lambda^{\prime}}{\sqrt{n-1}}\Big)\text{ and }(\sqrt{N-n},\sqrt{n-1})$
attains equality, which holds only if
$\frac{-\partial_{\gamma^{\prime}}\psi}{y^{\prime}-\lambda^{\prime}}=\frac{N-n}{n-1}.$
Thus, having $y^{\prime}-\lambda^{\prime}=\mathrm{tr}V=(n-1)\xi(t)$, we have $\partial_{\gamma^{\prime}}\psi=(N-n)\xi(t)$, and $\alpha^{\prime}=y^{\prime}-\lambda^{\prime}-\partial_{\gamma^{\prime}}\psi=(N-1)\xi(t)$. Hence equation (9) gives, when $\mathrm{Ric}_{m,N}=Kg$,
$\xi^{\prime}(t)+\xi^{2}(t)=-\frac{Kr^{2}}{N-1}\text{, where }r=d(x,\mathcal{F}_{1}(x))$
and the result on the sectional curvatures thus follows from the Riccati equation.
To derive information on the behaviour of measure along $\gamma(t)$, we recall the expression of the $N$-Ricci curvature
$\mathrm{Ric}_{m,N}=\mathrm{Ric}(\gamma^{\prime})+\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})-\frac{(\partial_{\gamma^{\prime}}\psi)^{2}}{N-n}.$
Given $\mathrm{Ric}_{m,N}(\gamma^{\prime})=Kr^{2}$ and $\mathrm{Ric}(\gamma^{\prime})=\frac{n-1}{N-1}Kr^{2}$, we can derive the following second order differential equation of $\psi$:
$\frac{d^{2}}{dt^{2}}\psi\circ\gamma(t)-\frac{1}{N-n}\Big[\frac{d}{dt}(\psi\circ\gamma(t))\Big]^{2}=\frac{N-n}{N-1}Kr^{2}.$
Therefore, it can be deduced that $e^{-\psi\circ\gamma(t)}$ satisfies the differential equation
$\frac{d^{2}}{dt^{2}}e^{\frac{-\psi\circ\gamma(t))}{N-n}}=-\frac{Kr^{2}}{N-1}e^{\frac{-\psi\circ\gamma(t)}{N-n}}.$
The solution of the differential equation yields
$e^{-\psi\circ\mathcal{F}_{t}(x)}=\left\{\begin{array}[]{cl}\Big[C_{0}\sin\Big(\sqrt{\frac{K}{N-1}}rt\Big)+C_{1}\cos\Big(\sqrt{\frac{K}{N-1}}rt\Big)\Big]^{N-n}\quad&\mbox{when $K>0$,}\\ (C_{0}t+C_{1})^{N-n}\quad&\mbox{when $K=0$,}\\ \Big[C_{0}\sinh\Big(\sqrt{\frac{-K}{N-1}}rt\Big)+C_{1}\cosh\Big(\sqrt{\frac{-K}{N-1}}rt\Big)\Big]^{N-n}\quad&\mbox{when $K<0$}\end{array}\right$.
where $C_{0}$ and $C_{1}$ are constants determined by the initial values $e^{-\psi(x)}$ and $\partial_{\gamma^{\prime}(0)}e^{-\psi(x)}$
$C_{0}=\frac{-\partial_{\gamma^{\prime}(0)}\psi(x)}{N-n}e^{\frac{-\psi(x)}{N-n}}\text{, }C_{1}=e^{\frac{-\psi(x)}{N-n}}.$
Remark 1.
When the dimension $n$ of $M$ is less than $N$, the density function $e^{-\psi\circ\mathcal{F}_{t}(x)}$ takes the value of 0 at $t=-\frac{1}{r}\sqrt{\frac{N-1}{K}}\arctan\big(\frac{C_{1}}{C_{0}}\big)$ when $K>0$; at $t=-\frac{C_{1}}{C_{0}}$ when $K=0$; or at $t=-\frac{1}{r}\sqrt{\frac{N-1}{-K}}\mathrm{arctanh}\big(\frac{C_{1}}{C_{0}}\big)$ when $K<0$ and $|C_{1}|<|C_{0}|$.
Nonetheless, this does not lead to a contradiction. On a smooth weighted Riemannian manifold $M$, the initial value $e^{-\psi(x)}$ is nonzero and the point at which the density $e^{-\psi}$ vanishes is not on the geodesic $\mathcal{F}_{t}(x)$ for all $t\in[0,1]$ for a.e. $x\in\mathrm{supp}f_{0}$, hence outside the support of $f_{0},f_{1}$ and $h$ and can be resolved by mollification. Therefore it is possible to attain equality of the Borell-Brascamp-Lieb inequality on smooth weighted Riemannian manifolds. (See also Remark 3.)
For completeness, we include also a rigidity result on unweighted Riemannian manifolds, which is a direct consequence of Theorem 2 and can be considered as an addition of the consideration of dimension bounds to the curvature rigidity result in Theorem 4.1 of [2].
Corollary 2.
On a Riemannian manifold $(M,g,\mathrm{vol}_{g})$ satisfying $\mathrm{Ric}_{N}\geq Kg$, rigidity of the Borell-Brascamp-Lieb inequality implies that the dimension $n$ of $M$ coincides with $N$ and the sectional curvatures along the geodesic $\gamma(t)$ are identically $\frac{K}{N-1}$.
4 The Brunn-Minkowski Inequality and its Rigidity
The Brunn-Minkowski inequality on Riemannian manifolds with $\mathrm{Ric}_{m,N}\geq Kg$, established in [20] is closely related to the Borell-Brascamp-Lieb inequality, whose rigidity can also be considered as a consequence of the rigidity of the Borell-Brascamp-Lieb inequality.
For Borel subsets $A,B\subset M$ of positive and finite measure, the Brunn-Minkowski inequality is a consequence of the Borell-Brascamp-Lieb inequality, by an appropriate selection of functions $f_{0},f_{1},h$ in the Borell-Brascamp-Lieb inequality (in the form of Corollary 1).
We pick $f_{0}:=\beta^{1-t}_{K,N}(\Theta_{A,B})\mathbf{1}_{A}$, $f_{1}:=\beta^{t}_{K,N}(\Theta_{A,B})\mathbf{1}_{B}$ and $h:=\mathbf{1}_{Z_{t}(A,B)}$ where $\Theta_{A,B}:=\inf_{(x,y)\in A\times B}d(x,y)$ when $K\geq 0$ and $\Theta_{A,B}:=\sup_{(x,y)\in A\times B}d(x,y)$ when $K<0$. It is easy to see that the condition $h(z)\geq\mathcal{M}^{p}_{t}\Big(\frac{f_{0}(x)}{\beta^{1-t}_{K,N}(d(x,y))},\frac{f_{1}(y)}{\beta^{t}_{K,N}(d(x,y))}\Big)$ for $z\in Z_{t}(x,y)$ is satisfied. Letting $p=+\infty$, the Borell-Brascamp-Lieb inequality implies the Brunn-Minkowski inequality:
$m(Z_{t}(A,B))=\int_{M}hdm\geq\mathcal{M}_{t}^{\frac{1}{N}}\Big(\int_{M}f_{0}dm,\int_{M}f_{1}dm\Big)\\ =\Big((1-t)\beta^{1-t}_{K,N}(\Theta_{A,B})^{\frac{1}{N}}m(A)^{\frac{1}{N}}+t\beta^{t}_{K,N}(\Theta_{A,B})^{\frac{1}{N}}m(B)^{\frac{1}{N}}\Big)^{N}.$
Remark 2.
The fact that on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$ the Borell-Brascamp-Lieb inequality implies Brunn-Minkowski inequality shows the equivalence between Borell-Brascamp-Lieb inequality and $\mathrm{Ric}_{m,N}\geq Kg$. $\mathrm{Ric}_{m,N}\geq Kg\Rightarrow$ BBL’ is known, and for the inverse implication, we have that the Borell-Brascamp-Lieb inequality implies Brunn-Minkowski inequality, which by [14] implies the $\mathrm{CD}(K,N)$ condition, equivalent to $\mathrm{Ric}_{m,N}\geq Kg$.
4.1 Rigidity in the null curvature bound
In the case of null-curvature bound, the Brunn-Minkowski inequality takes the form
$m(Z_{t}(A,B))^{\frac{1}{N}}\geq(1-t)m(A)^{\frac{1}{N}}+tm(B)^{\frac{1}{N}}.$
Assuming that the sets $A,B$ differ from convex sets up to a set of measure 0 and that $0<m(A),m(B)<\infty$, the result below is a rather straightforward consequence of the rigidity of the Borell-Brascamp-Lieb inequality in $\mathrm{CD}(0,N)$ manifolds.
Theorem 3.
Rigidity of the Brunn-Minkowski inequality on weighted Riemannian manifolds such that $\mathrm{Ric}_{m,N}\geq 0$ implies that for a.e. $x\in A$, the sectional curvatures along the geodesic of transport from $A$ to $B$ starting at $x$ are identically 0, and the weight behaves as a polynomial of degree NnN-n along the paths of transportation.
Proof.
Equality in the Brunn-Minkowski inequality reads
$m(Z_{t}(A,B))^{\frac{1}{N}}=(1-t)m(A)^{\frac{1}{N}}+tm(B)^{\frac{1}{N}}.$
Picking the functions to be the uniform distribution functions with supports being the respective sets, the result follows from the rigidity of the Borell-Brascamp-Lieb inequality, that for $m$-a.e. $x\in A$, along the geodesic $t\mapsto\mathcal{F}_{t}(x)$, the sectional curvatures are constantly 0 and the measure along it behaves as
$e^{-\psi\circ\mathcal{F}_{t}(x)}=(C_{0}t+C_{1})^{N-n}$
as in Theorem 2. ∎
Remark 3.
When $A$ is a point $\{x\}$ and $B$ is a ball centred at $x$, rigidity of the Brunn-Minkowski inequality would imply equality in the Bishop-Gromov volume comparison, which further implies that $B$ is a part of a flat cone due to the main result of [7], with a singularity at $\{x\}$ unless the dimension $n$ of $M$ is $N$. In general, equality in the Brunn-Minkowski inequality with $N>n$ does not necessarily imply the existence of a singularity, when $A$ and $B$ are set to be concentric annuli, the Brunn-Minkowski inequality about such $A$ and $B$ would be able to attain equality on a cone with its tip smoothened.
4.2 Rigidity in the positive and negative curvature bounds
The Brunn-Minkowski inequality takes the form
$m(Z_{t}(A,B))^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}\Big(\inf_{(x,y)\in A\times B}d(x,y)\Big)^{\frac{1}{N}}m(A)^{\frac{1}{N}}+t\beta^{t}_{K,N}\Big(\inf_{(x,y)\in A\times B}d(x,y)\Big)^{\frac{1}{N}}m(B)^{\frac{1}{N}}$
when $K>0$, and
$m(Z_{t}(A,B))^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}\Big(\sup_{(x,y)\in A\times B}d(x,y)\Big)^{\frac{1}{N}}m(A)^{\frac{1}{N}}+t\beta^{t}_{K,N}\Big(\sup_{(x,y)\in A\times B}d(x,y)\Big)^{\frac{1}{N}}m(B)^{\frac{1}{N}}$
when $K<0$.
By the scaling property of the $p$-mean, similar to the one used in the Theorem 4.1(ii) and Theorem 4.2 of [2], we have:
$1=\frac{\beta^{1-t}_{K,N}(\Theta_{A,B})}{\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))}=\frac{\beta^{t}_{K,N}(\Theta_{A,B})}{\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))}$ (11)
and due to the monotonicity of the function $\beta^{t}_{K,N}(r)$ in $r$, $d(x,\mathcal{F}_{1}(x))=\Theta_{A,B}$ for a.e. $x\in A$. Thus, by the same proof of Theorem 4.2 of [2], we conclude the following result on rigidity.
Theorem 4.
On a weighted Riemannian manifold with $\mathrm{Ric}_{m,N}\geq Kg$ where $K>0$, if compact convex sets $A,B\subset M$ with finite and positive measure satisfies equality of the Brunn-Minkowski inequality for some $t_{0}\in(0,1)$, then $A=B=Z_{t}(A,B)$ for all $t\in[0,1]$ up to a set of null measure. Equality of the Brunn-Minkowski inequality cannot hold for compact convex sets of finite and positive measure on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$ when $K<0$.
Proof.
For the case when $K>0$, we separate the situation when $A\cap B=\emptyset$ and when $A\cap B\neq\emptyset$.
When $A\cap B\neq\emptyset$, $\Theta_{A,B}=0$ and by equation (11), $\mathcal{F}_{1}(x)=x$ for almost all $x\in A$, implying that $A=B$ up to a set of measure 0, which further implies that $A=B=Z_{t}(A,B)$ up to a set of null measure.
When $A$ and $B$ are disjoint, we can apply the same argument constructed in [2] with $\frac{s_{K,N}(tx)}{s_{K,N}(x)}$ replaced by $\beta^{t}_{K,N}(x)$. We set $s:=\Theta_{A,B}=\inf_{(x,y)\in A\times B}d(x,y)=d(x,\mathcal{F}_{1}(x))$ for almost every $x\in A$. Fix $x\in A$ such that $d(x,\mathcal{F}_{1}(x))=s$, and $B_{s}(\mathcal{F}_{1}(x))\subset\bigcup_{y\in B}B_{s}(y)$. By construction, $B_{s}(\mathcal{F}_{1}(x))$ is disjoint with $A$. Fix $s_{0}\in(0,s)$ and for every $0<s^{\prime}<s_{0}$, for $z_{s^{\prime}}\in Z_{\frac{s^{\prime}}{2s}}(x,\mathcal{F}_{1}(x))$, $B_{\frac{s^{\prime}}{2}}(z_{s^{\prime}})$ is disjoint with $A$. This implies that $A$ is porous at $x$, that for every $s^{\prime}$ sufficiently small, there is a point $z_{s^{\prime}}$ of distance $s^{\prime}$ to $x$, such that there is an open ball with radius $\frac{s^{\prime}}{2}$ that is outside $A$ centred at which. In this case $m(A)$ cannot be positive, which is a contradiction to the assumption that $m(A)>0$.
In the case when $K<0$, by a similar construction of contradiction as in the case when $K>0$ and $A,B$ are disjoint, it can also be shown that for points $x\in A$ such that $d(x,\mathcal{F}_{1}(x))=\Theta_{A,B}$, $A$ is porous at $x$, implying that $m(A)=0$ which contradicts the assumption of $A$ having positive measure. ∎
5 Further Problems
Theorem 2 is concerned with the simplest situation of the rigidity of the Borell-Brascamp-Lieb inequality in metric measure geometry. There are a number of directions that it can be further generalised into.
  1. 1.
    After the result on rigidity of the Borell-Brascamp-Lieb inequality, it is natural to consider the quantitative stability of it on smooth weighted Riemannian manifolds with $\mathrm{Ric}_{m,N}\geq Kg$ for $N\in[n,\infty)$.
  2. 2.
    Another natural further problem is to consider the rigidity of the Borell-Brascamp-Lieb inequality with weaker dimension bounds, namely on smooth Riemannian manifolds with $\mathrm{Ric}_{m,N}\geq Kg$ for $N=\infty$ or $N\leq 0$.
  3. 3.
    It is possible to generalise Theorem 2 to Finsler manifolds, on which the curvature-dimension condition is described in [18], by considering the equality of the Finsler interpolation inequality presented in the same paper.
  4. 4.
    Following the proof of the Borell-Brascamp-Lieb inequality on metric measure spaces in [1], it is also natural to consider the rigidity of this inequality in a synthetic way, on $\mathrm{CD}(K,N)$ or $\mathrm{RCD}(K,N)$ metric measure spaces.
  5. 5.
    Considering the Borell-Brascamp-Lieb inequality together with its rigidity and stability on Lorentzian manifolds with timelike Ricci curvature bounds is also a direction of generalisation, under the framework of optimal transport on Lorentzian manifolds constructed in [16] and [17]. Recently, in [5], Cavalletti-Mondino proved the Brunn-Minkowski inequality on Lorentzian manifolds with timelike Ricci curvature bound, and in [9] Farooqui showed the equivalence between timelike Brunn-Minkowski inequality and timelike Ricci curvature bound.