### 20 Apr 2026 v2

# Rigidity of the Borell-Brascamp-Lieb Inequality on Weighted Riemannian Manifolds

Abstract
In this paper we discuss some results regarding the rigidity of the Borell-Brascamp-Lieb inequality and the Brunn-Minkowski inequality. We show a theorem of rigidity on curvature and measure of the Borell-Brascamp-Lieb inequality, a generalisation of the curvature rigidity theorem by Balogh-Kristály in (Balogh and Kristály, 2018) to the weighted setting. We present some rigidity results of the Brunn-Minkowski inequality and a few further open problems.
1 Introduction
Since the seminal works on the application of optimal transport in geometric analysis including (Otto and Villani, 2000), (Sturm, 2006) and (Lott and Villani, 2009), it became possible to study a wide range of analytical problems on Riemannian manifolds using the tool of optimal transport. In particular, the theory of optimal transport has been a powerful tool in the study of functional and geometrical inequalities on Riemannian manifolds with curvature and dimension bounds.
First proved in 1952 by Henstock and MacBeth on $\mathbb{R}$ by Henstock and Macbeath (1953) and later developed by Borell (1975) and Brascamp and Lieb (1976), the Borell-Brascamp-Lieb inequality is an integral inequality concerned with the $L^{1}$-norm of two functions and that of the mean of them.
On the $n$-dimensional Euclidean space $\mathbb{R}^{n}$, the Borell-Brascamp-Lieb inequality states that if functions $f,g,h>0$ such that for $s\in(0,1)$,
$h((1-s)x+sy)\geq\mathcal{M}^{p}_{s}(f(x),g(y))\quad\forall(x,y)\in\mathbb{R}^{n}\times\mathbb{R}^{n},$
then
$\|h\|_{L^{1}}\geq\mathcal{M}^{\frac{p}{1+np}}_{s}(\|f\|_{L^{1}},\|g\|_{L^{1}})$
where for $p\in[-\frac{1}{n},\infty]$ and real numbers $a,b\geq 0$,
[…]
is the $p$-mean, with the conventions $\mathcal{M}^{\infty}_{s}(a,b)=\max\{a,b\}$, $\mathcal{M}^{0}_{s}(a,b)=a^{1-s}b^{s}$ and $\mathcal{M}^{-\infty}_{s}=\min\{a,b\}$.
The Borell-Brascamp-Lieb inequality was proved on Riemannian manifolds by Cordero-Erausquin, McCann and Schmuckenschläger in (Cordero-Erausquin et al., 2001). It was then generalised further to metric measure spaces by Bacher (2010). The Borell-Brascamp-Lieb inequality implies a number of functional and geometric inequalities including the Brunn-Minkowski inequality and the Prékopa-Leindler inequality, a diagram showing their relationship can be seen in (Gardner, 2002).
Rigidity of the Borell-Brascamp-Lieb inequality on $\mathbb{R}^{n}$ has been studied since 1970s, various papers including (Dubuc, 1977) since then have discussed this problem. Rigidity of it on (unweighted) Riemannian manifolds with curvature bounded from below was shown by Balogh-Kristály in (Balogh and Kristály, 2018).
This paper generalises the curvature rigidity result of the Borell-Brascamp-Lieb inequality on Riemannian manifolds by Balogh and Kristály, Theorem 4.1 of (Balogh and Kristály, 2018), to weighted Riemannian manifolds. The main result of this paper states that the rigidity of the Borell-Brascamp-Lieb inequality implies identical sectional curvatures and certain behaviour of the measure.
Main Theorem (Theorem 2).
On a weighted Riemannian manifold $(M,g,m=\mathrm{e}^{-\psi}\mathrm{vol}_{g})$, satisfying $\mathrm{Ric}_{m,N}\geq Kg$ where $n\leq N<\infty$ and $K\in\mathbb{R}$, equality of the Borell-Brascamp-Lieb inequality implies that the $N$-Ricci curvature along the geodesic of transport $\gamma(t):=\mathcal{F}_{t}(x)$ equals $K$ and the sectional curvatures along the geodesic $\gamma$ are identically $\frac{K}{N-1}$. The measure behaves as an (Nn)(N-n)-th power of a linear combination of trignometric functions along the geodesic $\gamma(t)$ when $K>0$; as an (Nn)(N-n)-degree polynomial of $t$ along $\gamma(t)$ when $K=0$; and as an (Nn)(N-n)-th power of of a linear combination of hyperbolic functions along $\gamma(t)$ when $K<0$.
The article is structured as follows. In section 2, we first demonstrate a proof of an interpolation inequality on weighted Riemannian manifolds, followed by a proof of the Borell-Brascamp-Lieb inequality on weighted Riemannian manifolds. In section 3, we discuss the rigidity of the Borell-Brascamp-Lieb inequality and prove the main theorem, Theorem 2, and in section 4 we discuss the Brunn-Minkowski inequality and its rigidity, as a consequence of the rigidity of the Borell-Brascamp-Lieb inequality. At the end in section 5 we present a few related open questions.
Acknowledgements
I would like to express my deep gratitude to Prof. Shin-ichi Ohta for his guidance, suggestions and discussions. I would also like to thank Dr. Hiroshi Tsuji and Mr. Zisu Zhao for some helpful discussions.
This work was supported by JST SPRING, Grant Number JPMJSP2138.
2 Preliminaries
Throughout the paper, we consider a smooth weighted Riemannian manifold $(M,g,m)$ of dimension $n$, where $m=\mathrm{e}^{-\psi}\mathrm{vol}_{g}$, satisfying the condition $\mathrm{Ric}_{m,N}(\cdot)\geq Kg(\cdot,\cdot)$, $N\in[n,\infty)$ and $K\in\mathbb{R}$. Recall that the $N$-Ricci curvature $\mathrm{Ric}_{m,N}$ is defined to be:
$\mathrm{Ric}_{m,N}(v):=\left\{\begin{array}[]{cl}\mathrm{Ric}_{m}(v)-\frac{(\partial_{v}\psi)^{2}}{N-n}\quad&\mbox{for $N\in(n,\infty)$},\\ \mathrm{Ric}_{m}(v)\quad&\mbox{for $N=\infty$},\\ \mathrm{Ric}(v)\quad&\mbox{for $N=n$, $\psi=\mathrm{const.}$}\end{array}\right$.
where $\mathrm{Ric}_{m}(v)=\mathrm{Ric}(v)+\mathrm{Hess}_{\psi}(v,v)$.
We denote $\mathcal{P}(M)$ to be the space of Borel probability measures on $M$, and let $\mathcal{P}^{2}(M)$ be the space of measures with finite second moment
$\mathcal{P}^{2}(M)=\Big\{\mu\in\mathcal{P}(M)\Big|\int_{M}d(x,x_{0})^{2}d\mu<\infty\text{ for arbitrary }x_{0}\in M\Big\}.$
By the theorem of Brenier-McCann, given two measures on $(M,g,m)$, $\mu_{0}=\rho_{0}m=\rho_{0}\mathrm{e}^{-\psi}\mathrm{vol}_{g}$ and $\mu_{1}=\rho_{1}m=\rho_{1}\mathrm{e}^{-\psi}\mathrm{vol}_{g}$ in $\mathcal{P}^{2}(M)$ that are absolutely continuous with respect to $m$, there is a unique optimal transference plan from $\mu_{0}$ to $\mu_{1}$, given by $\mathcal{F}_{t}$, $\mu_{t}=(\mathcal{F}_{t})_{\#}\mu_{0}$, the push-forward of $\mu_{0}$ by $\mathcal{F}_{t}(x):=\exp_{x}(t\nabla f(x))$, and $f:M\rightarrow\mathbb{R}$ is a $\frac{d^{2}}{2}$-convex function. We refer to (Villani, 2003) for the fundamentals of optimal transport theory, see also (McCann, 2001) and (Figalli and Gigli, 2011) for the theory of optimal transport on Riemannian manifolds.
We consider the weighted Jacobian $J^{\psi}_{t}(x):=\mathrm{e}^{\psi(x)-\psi(\mathcal{F}_{t}(x))}\|(D\mathcal{F}_{t}(x))\|$ along the geodesic $\gamma(t):=\mathcal{F}_{t}(x)$. We have the following concavity condition on the weighted Jacobian:
$J^{\psi}_{t}(x)^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}$ (3)
for $\mu_{0}$-a.e. $x$ such that $\mathcal{F}_{t}$ is differentiable at $x$ and all $t\in[0,1]$, with the function $\beta^{t}_{K,N}$ being defined as:
$\beta^{t}_{K,N}(r)=\Big(\frac{s_{K,N}(tr)}{ts_{K,N}(r)}\Big)^{N-1}$
where
$s_{K,N}(r):=\left\{\begin{array}[]{cl}\sqrt{\frac{N-1}{K}}\sin\Big(r\sqrt{\frac{K}{N-1}}\Big)\quad&\mbox{if $K>0$},\\ r\quad&\mbox{if $K=0$},\\ \sqrt{-\frac{N-1}{K}}\sinh\Big(r\sqrt{\frac{-K}{N-1}}\Big)\quad&\mbox{if $K<0$}.\end{array}\right$.
It is known that the density function $\rho_{t}$ given by $\mu_{t}=\rho_{t}m$ satisfies the Monge-Ampère equation
$\rho_{0}(x)=\rho_{t}(\mathcal{F}_{t}(x))J^{\psi}_{t}(x)\text{ for }\mu_{0}\text{-a.e. }x.$ (5)
The interpolation inequality (3) on unweighted Riemannian manifolds was shown in the proof of Theorem 1.7 of (Sturm, 2006), and its Finsler version was shown in the proof of Theorem 1.2 of (Ohta, 2009). Here we present a proof on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$, along the computation shown in the proof of Theorem 1.7 of (Sturm, 2006), with an addition of the factor due to measure to the proof in (Sturm, 2006), for later use. See also pages 393-401 of (Villani, 2009).
Lemma 1.
On a weighted Riemannian manifold $(M,g,m=e^{-\psi(x)}\mathrm{vol}_{g})$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$, the weighted Jacobian $J^{\psi}_{t}(x):=e^{\psi(x)-\psi(\mathcal{F}_{t}(x))}\|(D\mathcal{F}_{t}(x))\|$ defined along the geodesic $\gamma(t):=\mathcal{F}_{t}(x)$, where $\mathcal{F}_{t}(x)=\exp_{x}(t\nabla f(x))$ described as above satisfies the concavity condition (3).
Proof.
We fix $x\in\mathrm{supp}\mu_{0}$ on $M$, and consider the geodesic
$\gamma(t):[0,1]\rightarrow M,t\mapsto\mathcal{F}_{t}(x)$
representing the path of transport starting from $x$.
When $\gamma(t)$ is constant, inequality (3) is trivial as $J_{t}^{\psi}(x)=1$ identically.
When $\gamma(t)$ is nonconstant, at $x$, we consider the tangent space $T_{x}M=T_{\mathcal{F}_{0}(x)}M$ and an orthonormal frame $\{e_{0}^{i}\}_{i=1}^{n}$, where we denote $e^{1}_{0}$ to be the unit tangent of the geodesic: $\frac{\gamma^{\prime}(0)}{\|\gamma^{\prime}(0)\|}$. We can thus define n1n-1 Jacobi fields $\mathbf{J}_{i}(t)=\frac{\partial}{\partial s}\exp_{x}(t(\nabla f(x)+se^{i}_{0}))$ for $i\in\{2,...,n\}$ that are orthogonal to $\gamma^{\prime}(t)$, and the vectors $\{\mathbf{J}_{i}(t)\}$ form a basis of $\{\gamma^{\prime}(t)\}^{\perp}\subset T_{\mathcal{F}_{t}(x)}M$ along the entire geodesic $\gamma$.
We denote $\mathrm{A}_{t}(x)$ to be the differential map:
$\mathrm{A}_{t}(x):=\mathrm{D}\mathcal{F}_{t}(x):T_{x}M\rightarrow T_{\mathcal{F}_{t}(x)}M.$
It is known that $\mathrm{A}_{t}$ satisfies the Riccati equation
$\mathrm{D}_{t}\mathrm{D}_{t}\mathrm{A}_{t}(x)+R(\mathrm{A}_{t}(x),\gamma^{\prime}(t))\gamma^{\prime}(t)=0$
with the initial condition $A_{0}=\mathrm{Id}$ and $\mathrm{D}_{t}A_{t}|_{t=0}=-\mathrm{Hess}\psi$, see page 142 of (Sturm, 2006).
Since $\{J_{i}\}_{i\geq 2}$ form a basis of the vector space $\{\gamma^{\prime}(t)\}^{\perp}$, and since for $i\geq 2$, $\langle\mathbf{J}_{i}^{\prime}(t),\gamma^{\prime}\rangle=0$ by Gauß lemma, we can define, as in (Sturm, 2006), the matrix $\mathbf{U}(t)$ to be such that $\mathbf{U}(t)\mathrm{A}_{t}:=\mathrm{D}_{t}\mathrm{A}_{t}$. And the Riccati equation can also be written as
$\mathbf{U}^{\prime}+\mathbf{U}^{2}+R\mathrm{A}_{t}^{-1}=0\text{, or }\mathbf{U}^{\prime}+\mathbf{U}^{2}+R(\cdot,\partial_{t}\mathcal{F}_{t})\partial_{t}\mathcal{F}_{t}=0.$
At $\gamma(t)$ we set an orthonormal basis $\{e^{i}_{t}\}$ with $e^{1}_{t}:=\gamma^{\prime}(t)/\|\gamma^{\prime}(t)\|$ being the tangent vector of $\gamma(t)$ of unit length. In terms of this basis, we set $\mathbf{U}_{ij}(t)=g(e^{i}_{t},\mathbf{U}(t)e^{j}_{t})$, $\lambda_{t}=1+\int_{0}^{t}U_{11}(s)ds$, and $f(t)=\exp(\lambda_{t})$, the Riccati equation implies that
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{U}_{11}(t)+\sum_{j=1}^{n}\mathbf{U}_{1j}^{2}=0\text{, therefore }-\lambda^{\prime\prime}=-\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{U}_{11}(t)=\sum_{j=1}^{n}\mathbf{U}_{1j}^{2}\geq\mathbf{U}_{11}^{2}=(\lambda^{\prime})^{2}$
which implies further that:
$f^{\prime\prime}=\lambda^{\prime\prime}f+(\lambda^{\prime})^{2}f\leq 0.$
We let $y_{t}:=\mathrm{log}J_{t}$ where $J_{t}:=\mathrm{det}\mathrm{A}_{t}=\|\mathrm{D}\mathcal{F}_{t}(x)\|$, $y^{\psi}_{t}:=\mathrm{log}J^{\psi}_{t}=\psi(x)-\psi\circ\mathcal{F}_{t}(x)+y_{t}$ and $\alpha_{t}:=y^{\psi}_{t}-\lambda_{t}$. Define
$u(t):=e^{\alpha_{t}}=\frac{J^{\psi}_{t}}{f(t)},$
since $y^{\prime}=\mathrm{tr}\mathbf{U}$, $y^{\prime\prime}=\mathrm{tr}(\mathbf{U}^{\prime})$, we have
$\alpha^{\prime\prime}=y^{\prime\prime}-\lambda_{t}^{\prime\prime}-(\psi\circ\mathcal{F}_{t})^{\prime\prime}$ $=$ $-\mathrm{Ric}(\gamma^{\prime})-\mathrm{tr}(\mathbf{U}^{2})-\mathbf{U}_{11}^{\prime}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$=$ $-\mathrm{Ric}(\gamma^{\prime})-\sum_{i,j=1}^{n}\mathbf{U}_{ij}^{2}+\sum_{j=1}^{n}\mathbf{U}_{1j}^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\sum_{i,j=2}^{n}\mathbf{U}_{ij}^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\frac{1}{n-1}\Big(\sum_{i=2}^{n}\mathbf{U}_{ii}\Big)^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime}).$
By the Cauchy-Schwarz inequality, it follows that
$\alpha^{\prime\prime}=y^{\prime\prime}-\lambda_{t}^{\prime\prime}-(\psi\circ\mathcal{F}_{t})^{\prime\prime}$ $\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\frac{(y^{\prime}-\lambda^{\prime})^{2}}{n-1}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$ (6)
$=$ $-\mathrm{Ric}_{m,N}(\gamma^{\prime})-\frac{(\partial_{\gamma^{\prime}}\psi)^{2}}{N-n}-\frac{(y^{\prime}-\lambda^{\prime})^{2}}{n-1}$
$\leq$ $-\mathrm{Ric}_{m,N}(\gamma^{\prime})-\frac{(\alpha^{\prime})^{2}}{N-1}.$
Inequality (6) and the hypothesis $\mathrm{Ric}_{m,N}\geq K$ imply that $u$ satisfies the differential inequality, as in pages 143-144 of (Sturm, 2006),
$(u^{\frac{1}{N-1}})^{\prime\prime}\leq-\frac{\mathrm{Ric}_{m,N}(\gamma^{\prime})}{N-1}u^{\frac{1}{N-1}}\leq-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}}\text{, where }r=d(x,\mathcal{F}_{1}(x))$
which implies that $u^{\frac{1}{N-1}}(t)$ satisfies the concavity condition
$u^{\frac{1}{N-1}}(t)\geq\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1).$
After multiplying $f^{\frac{1}{N}}$ and $u^{\frac{1}{N}}$ and applying the Hölder’s inequality, we have
$(J^{\psi}_{t})^{\frac{1}{N}}$ $=$ $(f(t)u(t))^{\frac{1}{N}}=f^{\frac{1}{N}}(t)(u^{\frac{1}{N-1}}(t))^{\frac{N-1}{N}}$
$\geq$ $((1-t)f(0)+tf(1))^{\frac{1}{N}}\Big(\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)\Big)^{\frac{N-1}{N}}$
$\geq$ $(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}$
showing the inequality (3). ∎
Along (Bacher, 2010), we can show a version of Borell-Brascamp-Lieb inequality on weighted Riemannian manifolds:
Theorem 1.
Let $f_{0},f_{1},h\geq 0$ be functions on a weighted Riemannian manifold $(M,g,m)$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$. Assume that for Borel subsets $A,B\subset M$, $\int_{A}f_{0}dm=\int_{B}f_{1}dm=1$ and
$h(z)^{-\frac{1}{N}}\leq(1-t)\beta^{1-t}_{K,N}(d(x,y))^{\frac{1}{N}}f_{0}(x)^{-\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,y))^{\frac{1}{N}}f_{1}(x)^{-\frac{1}{N}}$
for all $(x,y)\in A\times B,z\in Z_{t}(x,y)$, where
$Z_{t}(x,y):=\{z\in M|d(x,z)=td(x,y),d(z,y)=(1-t)d(x,y)\}$
is the set of $t$-intermediate points between $x$ and $y$. Then $\int_{M}hdm\geq 1$.
The proof of the above theorem is essentially the same as the proof of the main theorem of (Bacher, 2010), which is similar to the proof of the main theorem of (Cordero-Erausquin et al., 2001) by replacing the volume distortion coefficient by the function $\beta^{t}_{K,N}$.
Proof.
By the assumption that $\mathrm{Ric}_{m,N}\geq Kg$, we have the concavity condition of the weighted Jacobian (3).
We let $\rho_{0}:=f_{0}|_{A}$, $\rho_{1}:=f_{1}|_{B}$. There exists an optimal transference plan with density $\rho_{t}(x)$, satisfying Monge-Ampère equation (5), from $\mu_{0}=\rho_{0}m$ to $\mu_{1}=\rho_{1}m$ and induced by a map $\mathcal{F}_{t}$ such that
$J^{\psi}_{t}(x)\cdot\rho_{t}\circ\mathcal{F}_{t}(x)=\rho_{0}(x)=f_{0}(x)\text{ for }m\text{-a.e. }x\in A,$
and we have, letting $y=\mathcal{F}_{1}(x),$
$\Big(\frac{f_{0}(x)}{\rho_{t}(\mathcal{F}_{t}(x))}\Big)^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}(d(x,y))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,y))^{\frac{1}{N}}\Big(\frac{f_{0}(x)}{f_{1}(\mathcal{F}_{1}(x))}\Big)^{\frac{1}{N}},$
$\Big(\frac{1}{\rho_{t}(\mathcal{F}_{t}(x))}\Big)^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}(d(x,y))^{\frac{1}{N}}\frac{1}{f_{0}(x)^{\frac{1}{N}}}+t\beta^{t}_{K,N}(d(x,y))^{\frac{1}{N}}\frac{1}{f_{1}(\mathcal{F}_{1}(x))^{\frac{1}{N}}}\geq h(\mathcal{F}_{t}(x))^{-\frac{1}{N}},$
hence
$h\circ\mathcal{F}_{t}\geq\rho_{t}\circ\mathcal{F}_{t}.$
Thus, we conclude
$\int_{M}hdm$ $\geq$ $\int_{\mathcal{F}_{t}(A)}hdm\geq\int_{M}\mathbf{1}_{\mathcal{F}_{t}(A)}\rho_{t}dm$
$=\int_{A}\mathbf{1}_{\mathcal{F}_{t}(A)}\circ\mathcal{F}_{t}(x)\cdot\rho_{t}\circ\mathcal{F}_{t}(x)J^{\psi}_{t}dm=\int_{A}\mathbf{1}_{\mathcal{F}_{t}(A)}(\mathcal{F}_{t}(x))d\mu_{0}$
$=1.$
The above theorem can be expressed in terms of the $p$-mean via the same method of scaling as the main theorem of (Cordero-Erausquin et al., 2001) was generalised to Corollary 1.1 of (Cordero-Erausquin et al., 2001).
Corollary 1.
Let $f_{0},f_{1},h\geq 0$ be functions on a weighted Riemannian manifold $(M,g,m)$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$. Assume that
$h(z)\geq\mathcal{M}^{p}_{t}\Big(\frac{f_{0}(x)}{\beta^{1-t}_{K,N}(d(x,y))},\frac{f_{1}(y)}{\beta^{t}_{K,N}(d(x,y))}\Big)\text{ }\forall(x,y)\in M\times M,z\in Z_{t}(x,y)$
then
$\int_{M}hdm\geq\mathcal{M}_{t}^{\frac{p}{1+Np}}\Big(\int_{M}f_{0}dm,\int_{M}f_{1}dm\Big)$
where
[…]
is the $p$-mean, for $p\in\big[-\frac{1}{N},\infty\big]$, as introduced in the introduction.
3 Rigidity of the Borell-Brascamp-Lieb inequality
Suppose that for some $t_{0}\in(0,1)$, the Borell-Brascamp-Lieb inequality on a weighted Riemannian manifold as in Theorem 1 attains equality.
We therefore have, with the constructions above, for $\mu_{0}$-a.e. $x\in A$,
$\int_{M}hdm$ $=$ $\int_{\mathcal{F}_{t_{0}}(A)}hdm=\int_{M}\mathbf{1}_{\mathcal{F}_{t_{0}}(A)}\rho_{t_{0}}dm$
$=$ $\int_{A}\mathbf{1}_{\mathcal{F}_{t_{0}}(A)}\circ\mathcal{F}_{t_{0}}(x)\cdot\rho_{t_{0}}\circ\mathcal{F}_{t_{0}}(x)J^{\psi}_{t_{0}}dm=\int_{A}\mathbf{1}_{\mathcal{F}_{t_{0}}(A)}(\mathcal{F}_{t_{0}}(x))d\mu_{0}$
$=$ $1$
implying that on $A$, $h\circ\mathcal{F}_{t_{0}}=\rho_{t_{0}}\circ\mathcal{F}_{t_{0}}$ almost everywhere and hence
$J^{\psi}_{t_{0}}(x)^{\frac{1}{N}}=(1-t_{0})\beta^{1-t_{0}}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t_{0}\beta^{t_{0}}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}.$ (8)
By the comparison of differential inequalities, the interpolation inequality (8) attaining equality at one value of $t$ implies that the inequality attains equality for all $t\in[0,1]$, as follows.
Lemma 2.
If the weighted Jacobian $J^{\psi}_{t}$ satisfying (3) attains equality at some $t_{0}\in(0,1)$, then equality of the inequality (3) holds for all $t\in[0,1]$.
Proof.
The result is shown by comparison of ODE. We give a proof for $K>0$ and the other situations can be handled in the same way. Recall that with the definition of $f$ and $u$ as in the proof of Lemma 1, the inequality (3) is derived by:
$(J^{\psi}_{t})^{\frac{1}{N}}$ $=$ $(fu)^{\frac{1}{N}}=f^{\frac{1}{N}}(u^{\frac{1}{N-1}})^{\frac{N-1}{N}}$
$\geq$ $((1-t)f(0)+tf(1))^{\frac{1}{N}}\Big(\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)\Big)^{\frac{N-1}{N}}$
$\geq$ $(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}.$
Claim 1: If equality holds at $t_{0}$, then $f(t_{0})=(1-t_{0})f(0)+t_{0}f(1)$ and $u(t_{0})^{\frac{1}{N-1}}=\frac{s_{K,N}((1-t_{0})r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(t_{0}r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$ where $r=d(x,\mathcal{F}_{1}(x))$.
Proof of Claim 1: Suppose, on the contrary, that the claim is not true, that
$f(t_{0})>(1-t_{0})f(0)+t_{0}f(1)$
then we will have that $u(t_{0})^{\frac{1}{N-1}}<\frac{s_{K,N}((1-t_{0})r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(t_{0}r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$, which is a contradiction, as the differential inequality
$(u^{\frac{1}{N-1}})^{\prime\prime}\leq-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}}$
implies otherwise.
Or similarily, if $u(t_{0})^{\frac{1}{N-1}}>\frac{s_{K,N}((1-t_{0})r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(t_{0}r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$, we will get the same contradiction for $f$.
Claim 2: If $u$ satisfies $(u^{\frac{1}{N-1}})^{\prime\prime}\leq-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}}$ for $t\in(0,1)$ and there exists $t_{0}\in(0,1)$ such that $u(t_{0})^{\frac{1}{N-1}}=\frac{s_{K,N}((1-t_{0})r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(t_{0}r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$, then
$u^{\frac{1}{N-1}}(t)=\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)\quad\forall t\in[0,1].$
Proof of Claim 2: We define $w(t):=\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)$. Note that
$w^{\prime\prime}(t)=-\frac{Kr^{2}}{N-1}w.$
Thus, $u^{\frac{1}{N-1}}-w$ satisfies the differential inequality
$(u^{\frac{1}{N-1}}-w)^{\prime\prime}\leq\frac{Kr^{2}}{N-1}(w-u^{\frac{1}{N-1}})\leq 0$
and has boundary values $u^{\frac{1}{N-1}}(0)-w(0)=u^{\frac{1}{N-1}}(1)-w(1)=0$. Since also at $t_{0}$, $u^{\frac{1}{N-1}}-w=0$, by the maximum principle of subharmonic functions, we conclude
$u^{\frac{1}{N-1}}-w=0\quad\forall t\in[0,1].$
The lemma follows from Claim 2. ∎
We can thus show the main result by concluding:
Theorem 2.
Rigidity of the Borell-Brascamp-Lieb inequality on weighted Riemannian manifold $(M,g,m)$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$ implies, for a.e. $x\in\mathrm{supp}f_{0}$, letting $\gamma(t):=\mathcal{F}_{t}(x)$, the following hold:
  1. 1.
    For all $t\in(0,1)$, the $N$-Ricci curvature $\mathrm{Ric}_{m,N}(\gamma^{\prime}(t))$ equals $Kg(\gamma^{\prime},\gamma^{\prime})$, and the sectional curvatures along the geodesic $\gamma$ are identically $\frac{K}{N-1}$.
  2. 2.
    The measure behaves as an (Nn)(N-n)-th power of a linear combination of trignometric functions along the geodesic $\gamma(t)$ when $K>0$; as an (Nn)(N-n)-degree polynomial of $t$ along $\gamma(t)$ when $K=0$; and as an (Nn)(N-n)-th power of a linear combination of hyperbolic functions along $\gamma(t)$ when $K<0$.
Proof.
We assume $N>n$ in the proof, as the case of $N=n$ reduces to (Balogh and Kristály, 2018). With the above lemma and the construction of the functions $\alpha,y,\lambda_{t},u,f$ and the definition of the matrix $U$ as in the proof of Lemma 1, we have for all $t\in(0,1)$,
$(J^{\psi}_{t})^{\frac{1}{N}}$ $=$ $(fu)^{\frac{1}{N}}=f^{\frac{1}{N}}(u^{\frac{1}{N-1}})^{\frac{N-1}{N}}$
$=$ $((1-t)f(0)+tf(1))^{\frac{1}{N}}\Big(\frac{s_{K,N}((1-t)r)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(0)+\frac{s_{K,N}(tr)}{s_{K,N}(r)}u^{\frac{1}{N-1}}(1)\Big)^{\frac{N-1}{N}}$
$=$ $(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}$
which implies that f(t)=(1t)f(0)+tf(1)f(t)=(1-t)f(0)+tf(1) and the function $u$ satisfies
$(u^{\frac{1}{N-1}})^{\prime\prime}$ $=$ $-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}}.$
Since we have the chain of differential inequality
$(u^{\frac{1}{N-1}})^{\prime\prime}\leq-\frac{\mathrm{Ric}_{m,N}(\gamma^{\prime})}{N-1}u^{\frac{1}{N-1}}\leq-\frac{Kr^{2}}{N-1}u^{\frac{1}{N-1}},$
by squeezing, it can be concluded that along the geodesic $\gamma$, $\mathrm{Ric}_{m,N}(\gamma^{\prime})=Kg(\gamma^{\prime},\gamma^{\prime})$.
The differential inequality of $u$ attaining equality also implies that the differential inequality (6) attains equality:
$\alpha^{\prime\prime}=-\mathrm{Ric}_{m,N}(\gamma^{\prime})-\frac{(\alpha^{\prime})^{2}}{N-1}.$ (9)
The equality being attained has two implications. The first implication is that the matrix $V$, defined to be $\mathbf{V}:=(\mathbf{U}_{ij})_{i,j=2,...,n}$, satisfies the condition $\mathrm{tr}(\mathbf{V}^{2})=\frac{1}{n-1}(\mathrm{tr}\mathbf{V})^{2}$, that it equals the scalar multiple of the identity matrix $\mathbf{V}=\xi(t)\mathbf{\mathrm{Id}}_{n-1}$, which further implies that the sectional curvatures along the geodesic $\gamma(t)$ are all identical by the Riccati equation.
The second implication is that the Cauchy-Schwarz inequality between the tuples
$\Big(-\frac{\partial_{\gamma^{\prime}}\psi}{\sqrt{N-n}},\frac{y^{\prime}-\lambda^{\prime}}{\sqrt{n-1}}\Big)\text{ and }(\sqrt{N-n},\sqrt{n-1})$
attains equality, which holds only if
$\frac{-\partial_{\gamma^{\prime}}\psi}{y^{\prime}-\lambda^{\prime}}=\frac{N-n}{n-1}.$
Thus, having $y^{\prime}-\lambda^{\prime}=\mathrm{tr}\mathbf{V}=(n-1)\xi(t)$, we have $\partial_{\gamma^{\prime}}\psi=(N-n)\xi(t)$, and $\alpha^{\prime}=y^{\prime}-\lambda^{\prime}-\partial_{\gamma^{\prime}}\psi=(N-1)\xi(t)$. Hence equation (9) gives, when $\mathrm{Ric}_{m,N}=Kg$,
$\xi^{\prime}(t)+\xi^{2}(t)=-\frac{Kr^{2}}{N-1}\text{, where }r=d(x,\mathcal{F}_{1}(x))$
and the result on the sectional curvatures thus follows from the Riccati equation.
To derive information on the behaviour of measure along $\gamma(t)$, we recall the expression of the $N$-Ricci curvature
$\mathrm{Ric}_{m,N}=\mathrm{Ric}(\gamma^{\prime})+\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})-\frac{(\partial_{\gamma^{\prime}}\psi)^{2}}{N-n}.$
Given $\mathrm{Ric}_{m,N}(\gamma^{\prime})=Kr^{2}$ and $\mathrm{Ric}(\gamma^{\prime})=\frac{n-1}{N-1}Kr^{2}$, we can derive the following second order differential equation of $\psi$:
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\psi\circ\gamma(t)-\frac{1}{N-n}\Big[\frac{\mathrm{d}}{\mathrm{d}t}(\psi\circ\gamma(t))\Big]^{2}=\frac{N-n}{N-1}Kr^{2}.$
Therefore, it can be deduced that $\mathrm{e}^{-\psi\circ\gamma(t)}$ satisfies the differential equation
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathrm{e}^{\frac{-\psi\circ\gamma(t))}{N-n}}=-\frac{Kr^{2}}{N-1}\mathrm{e}^{\frac{-\psi\circ\gamma(t)}{N-n}}.$
The solution of the differential equation yields
$\mathrm{e}^{-\psi\circ\mathcal{F}_{t}(x)}=\left\{\begin{array}[]{cl}\Big[C_{0}\sin\Big(\sqrt{\frac{K}{N-1}}rt\Big)+C_{1}\cos\Big(\sqrt{\frac{K}{N-1}}rt\Big)\Big]^{N-n}\quad&\mbox{when $K>0$,}\\ (C_{0}t+C_{1})^{N-n}\quad&\mbox{when $K=0$,}\\ \Big[C_{0}\sinh\Big(\sqrt{\frac{-K}{N-1}}rt\Big)+C_{1}\cosh\Big(\sqrt{\frac{-K}{N-1}}rt\Big)\Big]^{N-n}\quad&\mbox{when $K<0$}\end{array}\right$.
where $C_{0}$ and $C_{1}$ are constants determined by the initial values $\mathrm{e}^{-\psi(x)}$ and $\partial_{\gamma^{\prime}(0)}\mathrm{e}^{-\psi(x)}$
$C_{0}=\frac{-\partial_{\gamma^{\prime}(0)}\psi(x)}{N-n}\mathrm{e}^{\frac{-\psi(x)}{N-n}}\text{, }C_{1}=\mathrm{e}^{\frac{-\psi(x)}{N-n}}.$
Remark 1.
When the dimension $n$ of $M$ is less than $N$, the density function $e^{-\psi\circ\mathcal{F}_{t}(x)}$ takes the value of 0 at $t=-\frac{1}{r}\sqrt{\frac{N-1}{K}}\arctan\big(\frac{C_{1}}{C_{0}}\big)$ when $K>0$; at $t=-\frac{C_{1}}{C_{0}}$ when $K=0$; or at $t=-\frac{1}{r}\sqrt{\frac{N-1}{-K}}\mathrm{arctanh}\big(\frac{C_{1}}{C_{0}}\big)$ when $K<0$ and $|C_{1}|<|C_{0}|$.
Nonetheless, this does not lead to a contradiction. On a smooth weighted Riemannian manifold $M$, the initial value $e^{-\psi(x)}$ is nonzero and the point at which the density $e^{-\psi}$ vanishes is not on the geodesic $\mathcal{F}_{t}(x)$ for all $t\in[0,1]$ for a.e. $x\in\mathrm{supp}f_{0}$, hence outside the support of $f_{0},f_{1}$ and $h$ and can be resolved by mollification. Therefore it is possible to attain equality of the Borell-Brascamp-Lieb inequality on smooth weighted Riemannian manifolds. (See also Remark 3.)
For completeness, we include also a rigidity result on unweighted Riemannian manifolds, which is a direct consequence of Theorem 2 and can be considered as an addition of the consideration of dimension bounds to the curvature rigidity result in Theorem 4.1 of (Balogh and Kristály, 2018).
Corollary 2.
On a Riemannian manifold $(M,g,\mathrm{vol}_{g})$ satisfying $\mathrm{Ric}_{N}\geq Kg$, rigidity of the Borell-Brascamp-Lieb inequality implies that the dimension $n$ of $M$ coincides with $N$ and the sectional curvatures along the geodesic $\gamma(t)$ are identically $\frac{K}{N-1}$.
4 The Brunn-Minkowski Inequality and its Rigidity
The Brunn-Minkowski inequality on Riemannian manifolds with $\mathrm{Ric}_{m,N}\geq Kg$, established in (Sturm, 2006) is closely related to the Borell-Brascamp-Lieb inequality, whose rigidity can also be considered as a consequence of the rigidity of the Borell-Brascamp-Lieb inequality.
For Borel subsets $A,B\subset M$ of positive and finite measure, the Brunn-Minkowski inequality is a consequence of the Borell-Brascamp-Lieb inequality, by an appropriate selection of functions $f_{0},f_{1},h$ in the Borell-Brascamp-Lieb inequality (in the form of Corollary 1).
We pick $f_{0}:=\beta^{1-t}_{K,N}(\Theta_{A,B})\mathbf{1}_{A}$, $f_{1}:=\beta^{t}_{K,N}(\Theta_{A,B})\mathbf{1}_{B}$ and $h:=\mathbf{1}_{Z_{t}(A,B)}$ where $\Theta_{A,B}:=\inf_{(x,y)\in A\times B}d(x,y)$ when $K\geq 0$ and $\Theta_{A,B}:=\sup_{(x,y)\in A\times B}d(x,y)$ when $K<0$. It is easy to see that the condition $h(z)\geq\mathcal{M}^{p}_{t}\Big(\frac{f_{0}(x)}{\beta^{1-t}_{K,N}(d(x,y))},\frac{f_{1}(y)}{\beta^{t}_{K,N}(d(x,y))}\Big)$ for $z\in Z_{t}(x,y)$ is satisfied. Letting $p=+\infty$, the Borell-Brascamp-Lieb inequality implies the Brunn-Minkowski inequality:
$m(Z_{t}(A,B))=\int_{M}hdm\geq\mathcal{M}_{t}^{\frac{1}{N}}\Big(\int_{M}f_{0}dm,\int_{M}f_{1}dm\Big)\\ =\Big((1-t)\beta^{1-t}_{K,N}(\Theta_{A,B})^{\frac{1}{N}}m(A)^{\frac{1}{N}}+t\beta^{t}_{K,N}(\Theta_{A,B})^{\frac{1}{N}}m(B)^{\frac{1}{N}}\Big)^{N}.$
Remark 2.
The fact that on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$ the Borell-Brascamp-Lieb inequality implies Brunn-Minkowski inequality shows the equivalence between Borell-Brascamp-Lieb inequality and $\mathrm{Ric}_{m,N}\geq Kg$. $\mathrm{Ric}_{m,N}\geq Kg\Rightarrow$ BBL’ is known, and for the inverse implication, we have that the Borell-Brascamp-Lieb inequality implies Brunn-Minkowski inequality, which by (Magnabosco et al., 2024) implies the $\mathrm{CD}(K,N)$ condition, equivalent to $\mathrm{Ric}_{m,N}\geq Kg$.
4.1 Rigidity in the null curvature bound
In the case of null-curvature bound, the Brunn-Minkowski inequality takes the form
$m(Z_{t}(A,B))^{\frac{1}{N}}\geq(1-t)m(A)^{\frac{1}{N}}+tm(B)^{\frac{1}{N}}.$ (11)
Assuming that the sets $A,B$ differ from convex sets up to a set of measure 0 and that $0<m(A),m(B)<\infty$, the result below is a rather straightforward consequence of the rigidity of the Borell-Brascamp-Lieb inequality in $\mathrm{CD}(0,N)$ manifolds.
Theorem 3.
Rigidity of the Brunn-Minkowski inequality on weighted Riemannian manifolds such that $\mathrm{Ric}_{m,N}\geq 0$ implies that for a.e. $x\in A$, the sectional curvatures along the geodesic of transport from $A$ to $B$ starting at $x$ are identically 0, and the weight behaves as a polynomial of degree NnN-n along the paths of transportation.
Proof.
Equality in the Brunn-Minkowski inequality reads
$m(Z_{t}(A,B))^{\frac{1}{N}}=(1-t)m(A)^{\frac{1}{N}}+tm(B)^{\frac{1}{N}}.$
Picking the functions to be the uniform distribution functions with supports being the respective sets, the result follows from the rigidity of the Borell-Brascamp-Lieb inequality, that for $m$-a.e. $x\in A$, along the geodesic $t\mapsto\mathcal{F}_{t}(x)$, the sectional curvatures are constantly 0 and the measure along it behaves as
$\mathrm{e}^{-\psi\circ\mathcal{F}_{t}(x)}=(C_{0}t+C_{1})^{N-n}$
as in Theorem 2. ∎
Remark 3.
When $A$ is a point $\{x\}$ and $B$ is a ball centred at $x$, rigidity of the Brunn-Minkowski inequality would imply equality in the Bishop-Gromov volume comparison, which further implies that $B$ is a part of a flat cone due to the main result of (De Philippis and Gigli, 2016), with a singularity at $\{x\}$ if $M$ is complete, unless the dimension $n$ of $M$ is $N$. Whilst it may be difficult to construct nontrivial examples of sets that attain equality of the Brunn-Minkowski inequality on complete smooth Riemannian manifolds with dimension less than $N$, it is possible to attain equality in the Brunn-Minkowski inequality on geodesically convex smooth open manifolds.
Example 1.
Let $M$ be the Euclidean half space $\mathbb{R}^{n}_{+}:=\{x=(x^{1},...,x^{n})\in\mathbb{R}^{n}|x^{n}>0\}$ endowed with the measure $m=\|x\|^{N-n}\mathcal{L}^{n}$. Let $A=B_{r}((0,...,1))$ be a ball of radius $r<1$ centred at (0,,1)(0,...,1) and $B=2A=B_{2r}((0,...,2))$. We have $Z_{t}(A,B)=(1+t)A$, and
$m(Z_{t}(A,B))$ $=$ $\int_{Z_{t}(A,B)}dm=\int_{(1+t)A}\|x\|^{N-n}d\mathcal{L}^{n}(x)=\int_{A}(1+t)^{n}\|(1+t)x\|^{N-n}d\mathcal{L}^{n}(x)$
$=$ $(1+t)^{N}m(A)=\Big((1-t)m(A)^{\frac{1}{N}}+tm(B)^{\frac{1}{N}}\Big)^{N}.$
4.2 Rigidity in the positive and negative curvature bounds
The Brunn-Minkowski inequality takes the form
$m(Z_{t}(A,B))^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}\Big(\inf_{(x,y)\in A\times B}d(x,y)\Big)^{\frac{1}{N}}m(A)^{\frac{1}{N}}+t\beta^{t}_{K,N}\Big(\inf_{(x,y)\in A\times B}d(x,y)\Big)^{\frac{1}{N}}m(B)^{\frac{1}{N}}$
when $K>0$, and
$m(Z_{t}(A,B))^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}\Big(\sup_{(x,y)\in A\times B}d(x,y)\Big)^{\frac{1}{N}}m(A)^{\frac{1}{N}}+t\beta^{t}_{K,N}\Big(\sup_{(x,y)\in A\times B}d(x,y)\Big)^{\frac{1}{N}}m(B)^{\frac{1}{N}}$
when $K<0$.
By the scaling property of the $p$-mean, similar to the one used in the Theorem 4.1(ii) and Theorem 4.2 of (Balogh and Kristály, 2018), we have:
$1=\frac{\beta^{1-t}_{K,N}(\Theta_{A,B})}{\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))}=\frac{\beta^{t}_{K,N}(\Theta_{A,B})}{\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))}$ (12)
and due to the monotonicity of the function $\beta^{t}_{K,N}(r)$ in $r$, $d(x,\mathcal{F}_{1}(x))=\Theta_{A,B}$ for a.e. $x\in A$. Thus, by the same proof of Theorem 4.2 of (Balogh and Kristály, 2018), we conclude the following result on rigidity.
Theorem 4.
On a weighted Riemannian manifold with $\mathrm{Ric}_{m,N}\geq Kg$ where $K>0$, if compact convex sets $A,B\subset M$ with finite and positive measure satisfies equality of the Brunn-Minkowski inequality for some $t_{0}\in(0,1)$, then $A=B=Z_{t}(A,B)$ for all $t\in[0,1]$ up to a set of null measure. Equality of the Brunn-Minkowski inequality cannot hold for compact convex sets of finite and positive measure on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$ when $K<0$.
Proof.
For the case when $K>0$, we separate the situation when $A\cap B=\emptyset$ and when $A\cap B\neq\emptyset$.
When $A\cap B\neq\emptyset$, $\Theta_{A,B}=0$ and by equation (12), $\mathcal{F}_{1}(x)=x$ for almost all $x\in A$, implying that $A=B$ up to a set of measure 0, which further implies that $A=B=Z_{t}(A,B)$ up to a set of null measure.
When $A$ and $B$ are disjoint, we can apply the same argument constructed in (Balogh and Kristály, 2018) with $\frac{s_{K,N}(tx)}{s_{K,N}(x)}$ replaced by $\beta^{t}_{K,N}(x)$. We set $s:=\Theta_{A,B}=\inf_{(x,y)\in A\times B}d(x,y)=d(x,\mathcal{F}_{1}(x))$ for almost every $x\in A$. Fix $x\in A$ such that $d(x,\mathcal{F}_{1}(x))=s$, and $B_{s}(\mathcal{F}_{1}(x))\subset\bigcup_{y\in B}B_{s}(y)$. By construction, $B_{s}(\mathcal{F}_{1}(x))$ is disjoint with $A$. Fix $s_{0}\in(0,s)$ and for every $0<s^{\prime}<s_{0}$, for $z_{s^{\prime}}\in Z_{\frac{s^{\prime}}{2s}}(x,\mathcal{F}_{1}(x))$, $B_{\frac{s^{\prime}}{2}}(z_{s^{\prime}})$ is disjoint with $A$. This implies that $A$ is porous at $x$, that for every $s^{\prime}$ sufficiently small, there is a point $z_{s^{\prime}}$ of distance $s^{\prime}$ to $x$, such that there is an open ball with radius $\frac{s^{\prime}}{2}$ that is outside $A$ centred at which. In this case $m(A)$ cannot be positive, which is a contradiction to the assumption that $m(A)>0$.
In the case when $K<0$, by a similar construction of contradiction as in the case when $K>0$ and $A,B$ are disjoint, it can also be shown that for points $x\in A$ such that $d(x,\mathcal{F}_{1}(x))=\Theta_{A,B}$, $A$ is porous at $x$, implying that $m(A)=0$ which contradicts the assumption of $A$ having positive measure. ∎
5 Further Problems
Theorem 2 is concerned with the simplest situation of the rigidity of the Borell-Brascamp-Lieb inequality in metric measure geometry. There are a number of directions that it can be further generalised into.
  1. 1.
    After the result on rigidity of the Borell-Brascamp-Lieb inequality, it is natural to consider the quantitative stability of it on smooth weighted Riemannian manifolds with $\mathrm{Ric}_{m,N}\geq Kg$ for $N\in[n,\infty)$.
  2. 2.
    Another natural further problem is to consider the rigidity of the Borell-Brascamp-Lieb inequality with weaker dimension bounds, namely on smooth Riemannian manifolds with $\mathrm{Ric}_{m,N}\geq Kg$ for $N=\infty$ or $N\leq 0$.
  3. 3.
    It is possible to generalise Theorem 2 to Finsler manifolds, on which the curvature-dimension condition is described in (Ohta, 2009), by considering the equality of the Finsler interpolation inequality presented in the same paper.
  4. 4.
    Following the proof of the Borell-Brascamp-Lieb inequality on metric measure spaces in (Bacher, 2010), it is also natural to consider the rigidity of this inequality in a synthetic way, on $\mathrm{CD}(K,N)$ or $\mathrm{RCD}(K,N)$ metric measure spaces.
  5. 5.
    Considering the Borell-Brascamp-Lieb inequality together with its rigidity and stability on Lorentzian manifolds with timelike Ricci curvature bounds is also a direction of generalisation, under the framework of optimal transport on Lorentzian manifolds constructed in (McCann, 2020) and (Mondino and Suhr, 2023). Recently, in (Cavalletti and Mondino, 2025), Cavalletti-Mondino proved the Brunn-Minkowski inequality on Lorentzian manifolds with timelike Ricci curvature bound, and in (Farooqui, 2025) Farooqui showed the equivalence between timelike Brunn-Minkowski inequality and timelike Ricci curvature bound.